# Enhancing Photometric Redshift Catalogs Through Color-Space Analysis: Application to KiDS-Bright Galaxies

### v1

@@ 3 Data
For our analysis, we combine photometric data originating from the VST and VISTA telescopes with spectroscopic redshifts taken from a number of surveys. Here, we provide details of the particular datasets.
@@ 3.1 Photometric data
Our photometric data is based primarily on the Kilo-Degree Survey (KiDS, de Jong et al., 2013) observations and associated data products. KiDS [^3] http://KiDS.strw.leidenuniv.nl obtained optical imaging in four bands ($ugri$) with the OmegaCAM instrument mounted at the Cassegrain focus of ESO’s VLT Survey Telescope (VST, Capaccioli & Schipani, 2011) at Cerro Paranal, Chile. These optical data are combined with their infrared $ZYJHK_{\rm s}$ counterparts from the VIsta Kilo degree INfrared Galaxy survey (VIKING, Edge et al., 2013), obtained using the Visible and InfraRed CAMera (VIRCAM) formerly mounted on ESO’s 4m VISTA telescope also on Cerro Paranal. The combined KiDS+VIKING dataset is very well matched in terms of depth and sky coverage (e.g. Wright et al., 2019).
@@
Here we use KiDS Data Release 4 (DR4, Kuijken et al., 2019), which covers roughly 1000 deg2 and includes 9-band photometric information. The source detection was performed on $r$-band imaging using the Source Extractor (Bertin & Arnouts, 1996) software, which also provides Kron-like AUTO magnitudes. Images are also post-processed to measure the Gaussian Aperture and PSF (GAaP, Kuijken, 2008) magnitudes. These were designed to yield accurate colors of galaxies (Kuijken et al., 2015) and hence are most appropriate for deriving color-sensitive quantities, such as photo-$z$ (see e.g., B18). We therefore use GAaP magnitudes also to build our feature space employed by the SOM. However, as discussed in e.g. Kuijken et al. (2019), GAaP magnitudes do not provide a good measurement of the total flux, especially for galaxies with large angular sizes, as in the case of KiDS-Bright. For the sample selection, we then follow B21 and use the $r$-band AUTO magnitudes to apply a flux limit. In what follows, we keep only those objects that have $r_{\mathrm{auto}}<20$. We also require 9-band GAaP measurements to be available. In order to obtain a galaxy (i.e. extended source) selection, this is accompanied by the following cuts [^4] See de Jong et al. (2015) for details of these morphological flags.: $\mathtt{CLASS\_STAR}<0.5$, $\mathtt{SG2DPHOT}=0$ and $\mathtt{SG\_FLAG}=1$. The KiDS catalog also provides mask information, which encodes (in particular) the areas affected by various artifacts. We follow the general recommendations for the DR4 selection and remove objects with $(\mathtt{MASK}\&28668)>0$. Finally, we remove from the KiDS-Bright sample any objects that were assigned $z_{\mathrm{phot}}<0$ (a rare artifact of the ML model, usually highlighting issues with photometry).
@@
For our analysis, we supplement the KiDS DR4 photometric data with optical+NIR measurements of similar quality as in KiDS+VIKING (KV) from the so-called ‘KiDZ’ fields. These are areas mostly external to the KiDS footprint, which were targeted with VST and VISTA observations both by independent surveys as well as via dedicated observations of the KiDS team. These fields overlap with various deep spectroscopic surveys, which allows the KiDS team to calibrate redshift distributions with direct approaches (such as DIR, Hildebrandt et al. 2017, 2020 or SOM, Wright et al. 2020b; Hildebrandt et al. 2021) thanks to having joint information of 9-band photometry and spectroscopy. For details of the KiDZ fields, see Wright et al. (2024). The data in the KiDZ fields include, for the vast majority of our sources, two separate measurements in the $i$-band to mimic the situation in KiDS DR5 (Wright et al., 2024). To homogenize these with the single $i$-band pass of DR4, we take the arithmetic mean of the magnitudes, which is a good approximation of mean flux for these bright, high signal-to-noise objects, especially that the two magnitudes are usually very consistent. If only one pass is available in KiDZ (a very rare occurrence, at a 0.02% level), we use it as the $i$-band measurement. We also implement masking and star removal, which here slightly differ from KiDS DR4, being in line with the DR5 post-processing.
@@
This study focuses on the KiDS-Bright sample (B21), a flux-limited galaxy dataset selected from KiDS DR4 as discussed above. The sample is accompanied by photo-$z$ generated with the neural network code ANNz2 (Sadeh et al., 2016), where the training set was derived from a cross-match between the KiDS and GAMA equatorial datasets. These photo-$z$ are based on 9-band GAaP magnitudes and show very good statistical accuracy and precision, namely mean bias (residual) $\langle\delta z\rangle=5\times 10^{-4}$ and scatter (scaled median absolute deviation from median, SMAD) $\sigma_{\delta z}=0.018(1+z)$. In what follows, we compare these photo-$z$ with spectroscopic redshifts via SOM projection.
@@
DESI EDR 0.261 19.82 12767 ZCAT_PRIMARY, $\mathtt{zWarn}=0$, $z>0.002$ DESI Collaboration et al. (2023)
@@
  • The table is sorted in descending number of sources within the KiDS DR4 + KiDZ bright sample coverage area. All the numbers apply after cross-matching the input spec-$z$ samples with the photometric data limited at $r<20$ mag and removing duplicates between surveys. We list only those surveys that have at least 100 cross-matched objects each.
In addition to the photo-$z$, in B21 a number of physical galaxy properties were derived for the KiDS-Bright dataset by employing the LePhare code (Arnouts et al., 1999; Ilbert et al., 2006) on the KV fluxes. Here we use one of these properties (galaxy stellar masses), which were shown in B21 to be consistent with the more accurate derivations from GAMA spectroscopy (e.g. Taylor et al., 2011). Stellar masses are relevant for our study, as these were employed by Dvornik et al. (2023) to bin the lens sample extracted from KiDS-Bright and to link the observed clustering and lensing signal of galaxies with the theoretical framework of the halo model (via the conditional stellar mass function).
@@ 3.2.1 GAMA
The basic spectroscopic dataset we use is derived from the Galaxy And Mass Assembly (GAMA, Driver et al., 2009) final Data Release 4 (Driver et al., 2022). The GAMA survey observed $\sim$ 300,000 galaxies down to $r\lesssim 19.8$ over 286 deg2 using the AAOmega multi-object spectrograph on the 4m Anglo-Australian Telescope. This includes five fields, of which four are fully contained within KiDS: three equatorial fields (G09, G12 and G15; Baldry et al., 2010), and one southern field (G23). The last GAMA field (G02, Liske et al., 2015) partly overlaps with KiDZ. In addition, a dedicated small-area, deeper survey lies within the G15 field (G15-Deep, Driver et al., 2022). Of these, the most spectroscopically complete are the equatorial fields (GAMA-eq hereafter), where redshift targets were originally selected from SDSS as flux-limited to $r_{\mathrm{Petro}}<19.8$ (with no color preselection) plus some additional filler targets fainter than this limit. The estimates of GAMA-eq completeness were originally at the level of $98.5\%$ for its fiducial flux limit (Liske et al., 2015). However, subsequent adoption of KiDS photometry for GAMA galaxies (Bellstedt et al., 2020) led to this being revised to $98\%$ at $r_{\mathrm{KiDS}}<19.6$ (Driver et al., 2022), where ‘KiDS’ refers to flux measurements made with the code ProFound (Robotham et al., 2018). (Bellstedt et al., 2020) then revised this completeness estimate to 95% at $r_{\mathrm{KiDS}}<19.72$ (Driver et al., 2022).
The GAMA-eq sample was used to calibrate the KiDS-Bright selection and photo-$z$ estimation in B21, and we will employ it as the main spectroscopic reference for our study. The other GAMA fields (G02 and G23; GAMA-noneq hereafter), have brighter flux limits and much less complete sampling than GAMA-eq (Liske et al., 2015), and will only serve as extensions together with other spec-$z$ samples discussed below. In all cases, GAMA redshifts are selected with the quality flag $\mathrm{NQ}\geq 3$ and with $z>0.002$ to avoid stellar contamination.
@@ 3.2.2 Extended spectroscopic compilation
In addition to the GAMA data detailed above, we use a number of other spectroscopic datasets overlapping with KiDS DR4 and KiDZ. Most of these are the same as used in KiDS cosmic shear studies and detailed in previous papers (Wright et al., 2024; van den Busch et al., 2022). However, we also add a few others and, most importantly, our cuts on the photometric data are different. The properties of these additional redshift samples are provided in Table 1. The numbers and statistics in the table apply after cross-matching with KiDS/KiDZ photometric data, cutting to $r_{\mathrm{auto}}<20$ to mimic the KiDS-Bright selection, and keeping only unique objects [^5] The matching is done within $1^{\prime\prime}$ radius. See Wright et al. (2024) for details on how duplicates are handled.. The datasets are listed in descending order of the number of sources within KiDS DR4 + KiDZ after the KiDS-Bright selections have been applied. We will denote this extended and matched spectroscopic compilation as ‘Espec’.
@@
Figure 2: The 30 × 30 hexagonal-cell SOMs with toroidal topology trained using the KiDS-Bright photometric sample (9 magnitudes+36 colors) of $\sim$1 million galaxies. These plots show a few example properties of the KiDS-Bright photometric sample projected on the SOM. From left to right: (1) $r$-band AUTO magnitude; (2) $g-r$ color; (3) logarithm of stellar mass in solar units; (4) photometric redshift derived with ANNz2. Properties (1) and (2) are taken from the KiDS DR4 photometric dataset (Kuijken et al., 2019), while (3) and (4) were obtained for the KiDS-Bright sample by B21. Colors indicate the mean value of a given property per SOM cell.
@@ 4 Color space analysis of KiDS-Bright galaxies
We train a $30\times 30$ hexagonal-cell SOM with toroidal topology using the KiDS-Bright photometric sample of $\sim$1 million galaxies. This gives us on average $\sim 1000$ galaxies per cell, which provides sufficient sampling per cell for our analysis. We train our SOM using nine magnitudes and 36 colors as our feature space and run the training for 100 epochs, which gives an optimal balance between acceptable map convergence and maintaining reasonable computational runtime. The learning scale is initialized with a value of 0.1 and is gradually reduced to 0.01 in the final epoch. The SOM, therefore, provides us with a 2D projection of the full 45 dimensions of KiDS-Bright color-magnitude space, which we utilize for our subsequent analyses. Furthermore, we also map various other parameters not used in the training of our SOM, including details of the spectroscopic samples discussed previously in Sec. 3.2.
Figure 3: Comparison of the number density of the KiDS-Bright photometric sample (left), GAMA spectroscopic (middle), both expressed as average sky-projected density per SOM cell and the ratio of the latter to the former (right), giving a measure of GAMA completeness with respect to KiDS-Bright. The ratio above unity can appear due to the cosmic variance between GAMA and KiDS datasets.
Figure 4: Dependence of the GAMA completeness (w.r.t. KiDS) on selected properties of the KiDS-Bright sample, based on SOM projections. The completeness is expressed as the per-cell ratio of GAMA and KiDS-Bright surface density. The panels show from left to right: $r$-band AUTO magnitude; $g-r$ color; logarithm of stellar mass; ANNz2 photometric redshifts. Each orange dot represents one SOM cell, while the blue points are median values with central 68-percentile error bars. Binning is done by preserving the same number of averaged cells per bin. The pink-shaded regions show the bootstrap error from 10000 bootstrap samples, as an approximation to the uncertainty on the median.
@@ 4.1 Comparison of KiDS-Bright and GAMA
We start by analyzing how various KiDS-Bright galaxy properties are mapped on the SOM. This allows us to make a qualitative comparison of how these quantities are correlated, including both the features employed to train the SOM and additional ones. In Fig. 2, we show four of them: $r$-band magnitude (first panel), $g-r$ color (second panel), stellar mass (third panel), and photometric redshift (fourth panel). The two former are taken directly from the observations (Kuijken et al., 2019); the two latter were derived from KiDS+VIKING photometry as described in B21. Stellar masses ($M_{*}$) were obtained using LePhare (Arnouts et al., 1999; Ilbert et al., 2006), and photo-$z$ were estimated using ANNz2 (Sadeh et al., 2016). Finally, we hereafter refer to $r$-band AUTO magnitudes simply as $r$ magnitudes, while colors are based on GAaP photometry.
@@
We start by noting that the projections from Fig. 2 look very similar between KiDS-Bright and GAMA, so we are not showing them here for the latter sample. However, more striking differences start to appear if we move to other quantities. An example is the number density ($\mathcal{N}$), which we obtain by dividing the galaxy counts of each SOM cell by the https://mocrefund.crcs.gov.in/effective survey area (777 deg2 for KiDS-Bright and 142 deg2 for GAMA). The result is plotted in Fig. 3 (left and middle panels). By comparing these with the left-hand side panel of Fig. 2, we see that the number density in the GAMA survey is much lower than in KiDS-Bright at the faint end of the latter. This is expected because the adopted KiDS-Bright flux limit of $r_{\mathrm{auto}}<20$ mag was slightly fainter than the GAMA-eq completeness limit of $r_{\mathrm{petro}}\lesssim 19.8$.
The right-hand panel of Fig. 3 illustrates the per-cell completeness defined as $\mathcal{C}\equiv\mathcal{N}_{\mathrm{GAMA}}$/$\mathcal{N}_{\mathrm{KiDS}}$. Here, completeness implies the parametric completeness of GAMA with respect to KiDS in magnitude. We note that this can be occasionally above unity, as for a given SOM training, the number of GAMA galaxies in some cells can be higher than that number for KiDS. This is due to two reasons: GAMA has more representative galaxies in a given cell, and also due to cosmic variance between GAMA and KiDS datasets. The faint-end incompleteness of GAMA, visible in the center of the middle and right panels of Fig. 3, does not need to mean that the KiDS-Bright photo-$z$ derived in B21 are unreliable there, provided that the GAMA selection at $r\sim 20$ mag seems to sample the KiDS-Bright color space sparsely (Fig. 2).
@@
The completeness $\mathcal{C}$, measured from the right-hand panel of Fig. 3, can be directly mapped to other properties through use of the SOM. Examples are provided in Fig. 4, where we show the completeness as a function of the same quantities as visualized in Fig. 2. Each individual point represents one SOM cell, and we see considerable scatter among these. This scatter is the result of a combination of sample variance, shot noise, and selection effects in the two samples. However, by averaging over the individual cells in bins of the respective $x$-axes, we can trace the overall patterns in the (in)completeness of GAMA-eq vs. KiDS-Bright. We illustrate this with the blue points indicating medians with the error bars reflecting the $(16,84)^{\rm th}$ percentile range.

### v2

@@ 3 Data
For our analysis, we combined photometric data originating from the VST and VISTA telescopes with spectroscopic redshifts taken from a number of surveys. Here, we provide details of the particular datasets.
@@ 3.1 Photometric data
Our photometric data are based primarily on the Kilo-Degree Survey (KiDS, de Jong et al., 2013) observations and associated data products. KiDS [^3] http://KiDS.strw.leidenuniv.nl obtained optical imaging in four bands ($ugri$) with the OmegaCAM instrument mounted at the Cassegrain focus of ESO’s VLT Survey Telescope (VST, Capaccioli & Schipani, 2011) at Cerro Paranal, Chile. These optical data are combined with their infrared $ZYJHK_{\rm s}$ counterparts from the VIsta Kilo degree INfrared Galaxy survey (VIKING, Edge et al., 2013), obtained using the Visible and InfraRed CAMera (VIRCAM) formerly mounted on ESO’s 4m VISTA telescope also on Cerro Paranal. The combined KiDS+VIKING dataset is very well matched in terms of the depth and sky coverage (e.g., Wright et al., 2019).
@@
Here we use KiDS Data Release 4 (DR4, Kuijken et al., 2019), which covers roughly 1000 deg2 and includes nine-band photometric information. The source detection was performed on $r$-band imaging using the Source Extractor (Bertin & Arnouts, 1996) software, which also provides Kron-like AUTO magnitudes. Images are also post-processed to measure the Gaussian aperture and PSF (GAaP, Kuijken, 2008) magnitudes. These were designed to yield accurate colors of galaxies (Kuijken et al., 2015) and, hence, they are most appropriate for deriving color-sensitive quantities, such as photo-$z$ (see e.g., B18). Therefore, we used GAaP magnitudes also to build our feature space employed by the SOM. However, as discussed in Kuijken et al. (2019), for instance, GAaP magnitudes do not provide a good measurement of the total flux, especially for galaxies with large angular sizes, as in the case of KiDS-Bright. For the sample selection, we then followed B21 and used the $r$-band AUTO magnitudes to apply a flux limit. In the following, we have retained only the objects that have $r_{\mathrm{auto}}<20$. We also require nine-band GAaP measurements to be available. To obtain a galaxy (i.e., extended source) selection, this is accompanied by the following cuts [^4] See de Jong et al. (2015) for details of these morphological flags.: $\mathtt{CLASS\_STAR}<0.5$, $\mathtt{SG2DPHOT}=0$ and $\mathtt{SG\_FLAG}=1$. The KiDS catalog also provides mask information, which encodes (in particular) the areas affected by various artifacts. We followed the general recommendations for the DR4 selection and removed objects with $(\mathtt{MASK}\&28668)>0$. Finally, from the KiDS-Bright sample, we removed any objects that were assigned $z_{\mathrm{phot}}<0$ (a rare artifact of the ML model, usually highlighting issues with photometry).
@@
For our analysis, we supplemented the KiDS DR4 photometric data with optical+NIR measurements of similar quality as in KiDS+VIKING (KV) from the so-called ”KiDZ” fields. These are areas mostly external to the KiDS footprint, which were targeted with VST and VISTA observations both by independent surveys and via dedicated observations of the KiDS team. These fields overlap with various deep spectroscopic surveys, which allows the KiDS team to calibrate redshift distributions with direct approaches (such as DIR, Hildebrandt et al. 2017, 2020, or SOM, Wright et al. 2020b; Hildebrandt et al. 2021) thanks to their having joint information of nine-band photometry and spectroscopy. For details of the KiDZ fields, see Wright et al. (2024). The data in the KiDZ fields include, for the vast majority of our sources, two separate measurements in the $i$-band to mimic the situation in KiDS DR5 (Wright et al., 2024). To homogenize these with the single $i$-band pass of DR4, we take the arithmetic mean of the magnitudes, which is a good approximation of mean flux for these bright objects with a high signal-to-noise ratio, especially given the two magnitudes are usually very consistent. If only one pass is available in KiDZ (a very rare occurrence, at a 0.02% level), we used it as the $i$-band measurement. We also implemented masking and star removal, which slightly differ from KiDS DR4 and are more in line with the DR5 post-processing.
@@
This study focuses on the KiDS-Bright sample (B21), a flux-limited galaxy dataset selected from KiDS DR4 as discussed above. The sample is accompanied by photo-$z$ generated with the neural network code ANNz2 (Sadeh et al., 2016), where the training set was derived from a cross-match between the KiDS and GAMA equatorial datasets. These photo-$z$ are based on nine-band GAaP magnitudes and show very good statistical accuracy and precision, namely, mean bias (residual) $\langle\delta z\rangle=5\times 10^{-4}$ and scatter (scaled median absolute deviation from median, SMAD) $\sigma_{\delta z}=0.018(1+z)$. In the following, we compare these photo-$z$ with spectroscopic redshifts via a SOM projection.
@@
DESI EDR 0.261 19.82 12767 ZCAT_PRIMARY, $\mathtt{zWarn}=0$, $z>0.002$ Collaboration et al. (2024)
@@
  • The table is arranged based on the descending number of sources within the KiDS DR4 + KiDZ bright sample coverage area. All the numbers apply after cross-matching the input spec-$z$ samples with the photometric data limited at $r<20$ mag and removing duplicates between surveys. We list only those surveys that have at least 100 cross-matched objects each.
In addition to the photo-$z$, in B21 a number of physical galaxy properties were derived for the KiDS-Bright dataset by employing the LePhare code (Arnouts et al., 1999; Ilbert et al., 2006) on the KV fluxes. Here, we use one of these properties (galaxy stellar masses), which were shown in B21 to be consistent with the more accurate derivations from GAMA spectroscopy (e.g. Taylor et al., 2011). Stellar masses are relevant for our study, as these were employed by Dvornik et al. (2023) to bin the lens sample extracted from KiDS-Bright and to link the observed clustering and lensing signal of galaxies with the theoretical framework of the halo model (via the conditional stellar mass function).
@@ 3.2.1 GAMA
The basic spectroscopic dataset we use is derived from the Galaxy And Mass Assembly (GAMA, Driver et al., 2009) final Data Release 4 (Driver et al., 2022). The GAMA survey observed $\sim$ 300,000 galaxies down to $r\lesssim 19.8$ over 286 deg2 using the AAOmega multi-object spectrograph on the 4m Anglo-Australian Telescope. This includes five fields, of which four are fully contained within KiDS: three equatorial fields (G09, G12 and G15; Baldry et al., 2010), and one southern field (G23). The last GAMA field (G02, Liske et al., 2015) partly overlaps with KiDZ. In addition, a dedicated small-area, deeper survey lies within the G15 field (G15-Deep, Driver et al., 2022). Of these, the most spectroscopically complete are the equatorial fields (GAMA-eq hereafter), where redshift targets were originally selected from SDSS as flux-limited to $r_{\mathrm{Petro}}<19.8$ (with no color preselection) plus some additional filler targets fainter than this limit. The estimates of GAMA-eq completeness were originally at the level of $98.5\%$ for its fiducial flux limit (Liske et al., 2015). However, subsequent adoption of KiDS photometry for GAMA galaxies (Bellstedt et al., 2020) led to this being revised to $98\%$ at $r_{\mathrm{KiDS}}<19.6$ (Driver et al., 2022), where ‘KiDS’ refers to flux measurements made with the code ProFound (Robotham et al., 2018). Bellstedt et al. (2020) then revised this completeness estimate to 95% at $r_{\mathrm{KiDS}}<19.72$ (Driver et al., 2022).
The GAMA-eq sample was used to calibrate the KiDS-Bright selection and photo-$z$ estimation in B21, and we will employ it as the main spectroscopic reference for our study. The other GAMA fields (G02 and G23; GAMA-noneq hereafter), have brighter flux limits and much less complete sampling than GAMA-eq (Liske et al., 2015), and only serve as extensions together with other spec-$z$ samples discussed below. In all cases, GAMA redshifts were selected with the quality flag $\mathrm{NQ}\geq 3$ and with $z>0.002$ to avoid stellar contamination.
@@ 3.2.2 Extended spectroscopic compilation
In addition to the GAMA data detailed above, we used a number of other spectroscopic datasets overlapping with KiDS DR4 and KiDZ. Most of these are the same as used in KiDS cosmic shear studies and detailed in previous papers (Wright et al., 2024; van den Busch et al., 2022). However, we also added a few others. Also, most importantly, our cuts on the photometric data are different. The properties of these additional redshift samples are provided in Table 1. The numbers and statistics in the table are applicable after a cross-matching with KiDS/KiDZ photometric data, cutting down to $r_{\mathrm{auto}}<20$ to mimic the KiDS-Bright selection and keeping only unique objects [^5] The matching is done within $1^{\prime\prime}$ radius. See Wright et al. (2024) for details on how duplicates are handled.. The datasets are listed in descending order of the number of sources within KiDS DR4 + KiDZ after the KiDS-Bright selections have been applied. We denote this extended and matched spectroscopic compilation as ‘Espec’.
@@ 4 Color space analysis of KiDS-Bright galaxies
We trained a $30\times 30$ hexagonal-cell SOM with toroidal topology using the KiDS-Bright photometric sample of $\sim$1 million galaxies. This gave us on average $\sim 1000$ galaxies per cell, which provides sufficient sampling per cell for our analysis. We trained our SOM using nine magnitudes and 36 colors as our feature space and run the training for 100 epochs, which gave an optimal balance between acceptable map convergence and maintaining reasonable computational runtime. The learning scale was initialized with a value of 0.1 and is gradually reduced to 0.01 in the final epoch. The SOM, therefore, provides us with a 2D projection of the full 45 dimensions of KiDS-Bright color-magnitude space, which we utilized for our subsequent analyses. Furthermore, we also mapped various other parameters not used in the training of our SOM, including details of the spectroscopic samples discussed previously in Sect. 3.2.
Figure 2: 30 $\times$ 30 hexagonal-cell SOMs with toroidal topology trained using the KiDS-Bright photometric sample (9 magnitudes+36 colors) of $\sim$1 million galaxies. These plots show a few example properties of the KiDS-Bright photometric sample projected on the SOM. From left to right: (1) $r$-band AUTO magnitude, (2) $g-r$ color, (3) logarithm of stellar mass in solar units, and (4) photometric redshift derived with ANNz2. Properties (1) and (2) are taken from the KiDS DR4 photometric dataset (Kuijken et al., 2019), while (3) and (4) were obtained for the KiDS-Bright sample by B21. Colors indicate the mean value of a given property per SOM cell.
Figure 3: Comparison of the number density of the KiDS-Bright photometric sample (left), GAMA spectroscopic (middle), both expressed as average sky-projected density per SOM cell and the ratio of the latter to the former (right), giving a measure of GAMA completeness with respect to KiDS-Bright. The ratio above unity can appear as a result of the cosmic variance between GAMA and KiDS datasets.
Figure 4: Dependence of the GAMA completeness (w.r.t. KiDS) on selected properties of the KiDS-Bright sample, based on SOM projections. The completeness is expressed as the per-cell ratio of GAMA and KiDS-Bright surface density. The panels show from left to right: $r$-band AUTO magnitude, $g-r$ color, logarithm of stellar mass, and ANNz2 photometric redshifts. Each orange dot represents one SOM cell, while the blue points are median values with central 68-percentile error bars. Binning is done by preserving the same number of averaged cells per bin. The pink-shaded regions show the bootstrap error from 10000 bootstrap samples, as an approximation to the uncertainty on the median.
@@ 4.1 Comparison of KiDS-Bright and GAMA
We start by analyzing how various KiDS-Bright galaxy properties are mapped on the SOM. This allows us to make a qualitative comparison of how these quantities are correlated, including both the features employed to train the SOM and additional ones. In Fig. 2, we show four of them: $r$-band magnitude (first panel), $g-r$ color (second panel), stellar mass (third panel), and photometric redshift (fourth panel). The two former are taken directly from the observations (Kuijken et al., 2019), the two latter were derived from KiDS+VIKING photometry as described in B21. Stellar masses ($M_{*}$) were obtained using LePhare (Arnouts et al., 1999; Ilbert et al., 2006), and photo-$z$ were estimated using ANNz2 (Sadeh et al., 2016). Finally, we hereafter refer to $r$-band AUTO magnitudes simply as $r$ magnitudes, while colors are based on GAaP photometry.
@@
We start by noting that the projections from Fig. 2 look very similar between KiDS-Bright and GAMA, so we are not showing them here for the latter sample. However, more striking differences start to appear if we move to other quantities. An example is the number density ($\mathcal{N}$), which we obtain by dividing the galaxy counts of each SOM cell by the effective survey area (777 deg2 for KiDS-Bright and 142 deg2 for GAMA). The result is plotted in Fig. 3 (left and middle panels). By comparing these with the left-hand side panel of Fig. 2, we see that the number density in the GAMA survey is much lower than in KiDS-Bright at the faint end of the latter. This is expected because the adopted KiDS-Bright flux limit of $r_{\mathrm{auto}}<20$ mag was slightly fainter than the GAMA-eq completeness limit of $r_{\mathrm{petro}}\lesssim 19.8$.
The right-hand panel of Fig. 3 illustrates the per-cell completeness defined as $\mathcal{C}\equiv\mathcal{N}_{\mathrm{GAMA}}$/$\mathcal{N}_{\mathrm{KiDS}}$. Here, completeness implies the parametric completeness of GAMA with respect to KiDS in magnitude. We note that this can be occasionally above unity, as for a given SOM training, the number of GAMA galaxies in some cells can be higher than that number for KiDS. This is due to two reasons: GAMA has more representative galaxies in a given cell, and also due to cosmic variance between GAMA and KiDS datasets. The faint-end incompleteness of GAMA, visible in the center of the middle and right panels of Fig. 3, does not need to mean that the KiDS-Bright photo-$z$ derived in B21 are unreliable there, provided that the GAMA selection at $r\sim 20$ mag seems to sparsely sample the KiDS-Bright color space (Fig. 2).
@@
The completeness, $\mathcal{C}$, measured from the right-hand panel of Fig. 3, can be directly mapped to other properties through the use of the SOM. Examples are provided in Fig. 4, where we show the completeness as a function of the same quantities as visualized in Fig. 2. Each individual point represents one SOM cell, and we see considerable scatter among these. This scatter is the result of a combination of sample variance, shot noise, and selection effects in the two samples. However, by averaging over the individual cells in bins of the respective $x$-axes, we can trace the overall patterns in the (in)completeness of GAMA-eq vs. KiDS-Bright. We illustrate this with the blue points indicating medians with the error bars reflecting the $(16,84)^{\rm th}$ percentile range.